# Bootstrapping World Models from Dynamics Models in Multimodal Foundation Models

### v1

@@ A.1 Limitations
While our approach demonstrates the effectiveness of our approaches across Aurora-Bench, the authors would like to highlight few limitations we have discovered:
@@
  • Despite efforts to guide the model via supervised fine-tuning with loss weighting or inference-time verification (Table 2), we observe that the model may still resort to copying the source observation, especially under low sampling temperatures or ambiguous instructions.
@@
  • While we show preliminary results of language-steered observation prediction in Figure 9, fine-grained control remains limited, and understanding subtle instructions (e.g., spatial or quantitative edits) remains challenging.
@@
  • We observe variance across different runs of experiments, likely due to the sensitivity of sampling for generation in multimodal models. To address this, we report results averaged over multiple runs and include performance under the best-of-$N$ sampling distribution during inference for a robust comparison.
  • We mostly conduct experiments using the native and unified VLM, Chameleon, as it is currently the only open-source VLM that supports interleaved image-text generation by default. This choice allows for fair and consistent benchmarking across our tasks. Moreover, Chameleon has demonstrated competitive performance in our settings. For example, its results are comparable to VILA-U in our dynamics prediction task. Future work should explore the generalisation to other multimodal foundation models with stronger capabilities.
A.2 Broader Impact
@@
A.3 Model Performance on Aurora-Bench with 5 Metrics
In addition to GPT4o-as-a-judge evaluation, we further employ a diverse set of automatic metrics covering both low-level and semantic fidelity: 1) we compute the L1 distance between the predicted and target observation as a pixel-level metric. 2) We extract visual features and compute the cosine similarity in their respective embedding spaces for several image encoders, including (CLIP-I and DINO), to assess semantic similarity. Additionally, to measure alignment between image content and the action semantics, we compute CLIP-T, the similarity between the edited image and its BLIP-generated caption. These metrics are evaluated in addition to GPT4o-as-a-judge metric following previous works in image editing [31, 30, 18]. We report the detailed results with 5 metrics in Table 3. We notice that copy baseline exhibits the best performance as measured by the distance-based and visual encoder-based approach, as indicated in Table 2. This poses a challenge to the reliability of the traditional metrics in fairly evaluating the action-centric image editing task. On the other hand, GPT4o-as-a-judge metric robustly assigns 00 score to Copy, indicating its robustness in detecting copying generation while putting GPT-as-a-judge as the most reliable metric to interpret.
Table 3: Model performance at MagicBrush, Action-Genome, Something, WhatsUp and Kubric on Aurora-Bench. For C-FT and CWM We report both the model performance and their performance in the best-of-N distribution. We report the average GPT4o scores for each model at the bottom. We highlight the better GPT-4o scores for C-FT and CWM. We bold the best performance among all models, except Copy and best-of-N performances. SE: SmartEdit.
@@ A.3 Model Performance on Aurora-Bench with 5 Metrics
Copy PixInstruct GoT SE CM C-FT +Best-of-3 CWM +Best-of-3
@@
A.4 Qualitative Cases
In this section, we present additional qualitative examples from Aurora-Bench in Figure 10. We observe several common failure modes in image editing models. First, they sometimes fail to preserve the scene from the source observation (e.g., PixInstruct on Action-Genome and MagicBrush). Second, some models generate near-identical copies of the source as the target (e.g., GoT on Something-Something). Third, producing realistic outputs remains difficult, as seen in GoT’s result on Kubric. Finally, maintaining object consistency is also a challenge—SmartEdit alters the object in WhatsUp, and CWM does so in Something-Something.
Despite the challenges, we also observe several positive editing behaviours from CWM. On Action-Genome, CWM correctly predicts spatial changes, such as opening and closing a drawer, which requires a strong understanding of the spatial concepts. In Something-Something, it is the only model to accurately capture the spatial concept of “falling down.” On Kubric, it demonstrates basic counting ability by correctly adding one keyboard. In WhatsUp, CWM correctly grounds the action to the laptop, while other models mistakenly edit the monitor.
@@ A.4 Qualitative Cases
Figure 10: Qualitative examples of the predicted next observation from the state-of-the-art specialised image editing models, and our models including C-FT and CWM, on Aurora-Bench.
A.5 Details of Processing CDM Annotations for Unlabelled Videos
@@
Figure 11: Distributions of triplet log-likelihoods predicted by CDM on Movements-in-Time, UCF-101, and Kinetics-700, based on 7K synthetic triplets per dataset. Triplets are uniformly sampled from each action class while maximising overall predicted likelihoods.
@@
[…]
@@
[…]
@@
We present the raw dataset statistics before sampling for Movements-in-Time, UCF-101 and Kinetics700 in Table 4. Figure 11 shows the distribution of CDM’s predicted scores across action classes in Movements-in-Time, Kinetics700, and UCF-101. The predicted likelihoods are nearly uniform within each class, indicating that our sampling method maintains both class diversity and high overall likelihoods. The sampling procedure for CDM-annotated trajectories is detailed in Algorithm 1.
Table 4: Dataset statistics for the video and triplets from the trajectories annotated by CDM. OPV: observations (i.e., extracted key-frames) per video, APV: actions per video, WPA: words per action.
@@
A.6 Prompt template for using GPT4o-as-a-Judge evaluation.
We provide the prompts used for evaluating image editing performance with GPT-4o in Figure A.6. We use GPT-4o-2024-11-20. The final score is the average of the minimum value of the two scores for each sample, as in [30].
Prompt Template for GPT4o-as-a-judge Evaluation You are a professional digital artist. You will have to evaluate the effectiveness of the AI-generated image(s) based on the given rules. You will have to give your output in a valid way of a Python dictionary format (Keep your reasoning concise and short.): {{"score": [...], "reasoning": "..." }}
and don’t output anything else. Two images will be provided: The first being the original AI-generated image The second being an edited version of the first. The objective is to evaluate how successfully the editing instruction has been executed in the second image. Note that sometimes the two images might look identical due to a failure in image editing. From a scale of 0 to 10: A score from 0 to 10 will be given based on the success of the editing. 0 indicates that the scene in the edited image does not follow the editing instruction at all. 10 indicates that the scene in the edited image follows the editing instruction text perfectly. If the object in the instruction is not present in the original image at all, the score will be 0. A second score from 0 to 10 will rate the degree of minimal editing in the second image. 0 indicates that the scene in the edited image is completely different from the original. 10 indicates that the edited image can be recognised as a minimally edited yet effective version of the original. Put the score in a list such that: output score = [score1, score2], where score1 evaluates the editing success and score2 evaluates the degree of the minimal editing. Editing instruction: {instruction}
A.7 Detailed GPT4o Scores for Editing Success and Minimal Editing
(a) Detailed GPT4o scores for CWM trained with the standard loss.
(b) Detailed GPT4o scores for CWM trained with the L2-weighted loss.
Figure 12: GPT4o scores’ distributions of editing success (ES) and minimal editing (OE) for CWM trained with standard loss or our loss-weighting method.
@@ A.7 Detailed GPT4o Scores for Editing Success and Minimal Editing
A.8 Implementation Details
A.8.1 Chameleon Dynamics Model
@@
We fine-tune the Chameleon-7B checkpoint from the Anole-7B version [51] to predict the action given a pair of observations, framed as an action-prediction task. The model is trained on a merged dataset from Action-Genome, Kubric, MagicBrush, Something-Something from AURORA’s annotated trajectories, and 15K EPIC-Kitchens processed by us. We downsample Kubric’s trajectories to 10K. Training is performed for 10 epochs with a batch size of 64, using a learning rate of 2e-4 and cosine scheduling (500 warm-up steps). We use bfloat16 mixed-precision training and apply LoRA [52] for parameter-efficient fine-tuning (rank 16, $\alpha=32$, dropout 0.05). Only the completion loss is used to optimise the generation of action. Training is conducted on 4 NVIDIA-H100-80GB-HBM3 GPUs using DeepSpeed for distributed optimisation.
A.8.2 C-FT Baseline
@@
We fine-tune the Chameleon-7B checkpoint from the Anole-7B version [51]. The model is trained on a combined dataset from Action-Genome, Kubric, MagicBrush, and Something-Something, formatted as the image editing task. We downsample Kubric’s trajectories to 10K. Training is conducted for 40 epochs with a batch size of 96 using the AdamW optimiser and a cosine learning rate scheduler (learning rate of 5e-4, 400 warm-up steps). We use mixed-precision training with bfloat16 and apply LoRA [52] for efficient fine-tuning (rank 16, $\alpha=32$, dropout 0.05). We only train the model with the truncated loss from the completion part. We use 4 NVIDIA-H100-80GB-HBM3 GPUs with DeepSpeed for distributed training. During inference, we apply a logits processor to mask out non-image tokens, set the temperature to 1, and use top-1 sampling. We observe that temperature is critical in controlling model behaviour: lower values often cause the model to copy the source observation instead of generating meaningful edits.
A.8.3 Chameleon World Model
@@
We fine-tune the Chameleon-7B checkpoint from the Anole-7B version [51]. The model is trained on a combined dataset from Action-Genome, Kubric, MagicBrush, Something-Something from AURORA’s annotated trajectories, together with 7K trajectories from Movements-in-Time, 7K trajectories from UCF-101 and 7K trajectories from Kinetics700, formatted as the image editing task. Again, we downsample Kubric’s trajectories to 10K. Training is conducted for 40 epochs with a batch size of 96 using the AdamW optimiser and a cosine learning rate scheduler (learning rate of 5e-4, 400 warm-up steps). We use mixed-precision training with bfloat16 and apply LoRA [52] for efficient fine-tuning (rank 16, $\alpha=32$, dropout 0.05). We only train the model with the truncated loss from the completion part, but we weight the image tokens using L2 strategy as introduced in Section 3. We use 4 NVIDIA-H100-80GB-HBM3 GPUs with DeepSpeed for distributed training. We use the same hyperparameters as C-FT during the inference time.
A.8.4 Computing Resources
@@
All training experiments were conducted on a compute node equipped with 4× NVIDIA H100 80GB GPUs, 256 CPU cores, and 256GB of memory. The total GPU hours required for training C-FT, CWM, and CDM were approximately 200, 400, and 100 hours, respectively.
@@
For inference, we used a single NVIDIA A100 80GB GPU with 8 CPU cores and 128GB memory. Inference for C-FT and CWM takes approximately 1 GPU hour per model. When applying verification with $K=8$, inference time increases to around 8 GPU hours. CDM only takes around 0.3 GPU hours for inference.
A.8.5 Assets and Licenses
@@
A.9 Details of Human Evaluation
@@
A.10 Safeguards
CWM performs observation prediction through image generation and, while its outputs are task-specific, we acknowledge that any generative model may carry potential for misuse. To mitigate these risks, we commit to the following safeguards upon release:

### v2

@@ 6 Conclusions
While overall our results demonstrate the effectiveness of our approaches across Aurora-Bench, we would like to highlight a few limitations that we have discovered:
@@
  • Despite efforts to guide the model via weakly supervised fine-tuning with loss weighting or inference-time verification (Table 2), we observe that the model may still resort to copying the source observation, especially under low sampling temperatures or ambiguous instructions.
@@
  • While we show promising preliminary results of language-steered next-observation prediction in Figure 5, fine-grained control remains limited, and understanding subtle instructions (e.g., spatial or quantitative edits) remains challenging.
@@
  • We observe high variance across different runs of experiments for Chameleon, likely due to the sensitivity of sampling from a weak model. To address this, we report results averaged over multiple runs.
Appendix A Potential Risks
@@
Appendix B Model Performance on Aurora-Bench with 5 Metrics
In addition to GPT4o-as-a-judge evaluation, we further employ a diverse set of automatic metrics covering both low-level and semantic fidelity: 1) we compute the L1 distance between the predicted and target observation as a pixel-level metric. 2) We extract visual features and compute the cosine similarity in their respective embedding spaces for several image encoders, including (CLIP-I and DINO), to assess semantic similarity. Additionally, to measure alignment between image content and the action semantics, we compute CLIP-T, the similarity between the edited image and its BLIP-generated caption. These metrics are evaluated in addition to GPT4o-as-a-judge metric following previous works in image editing (Huang et al., 2024; Fang et al., 2025; Krojer et al., 2024). We report the detailed results with 5 metrics in Table 6. We notice that copy baseline exhibits the best performance as measured by the distance-based and visual encoder-based approach, as indicated in Table 2. This poses a challenge to the reliability of the traditional metrics in fairly evaluating the action-centric image editing task. On the other hand, GPT4o-as-a-judge metric robustly assigns 0 score to Copy, indicating its robustness in detecting copying generation while putting GPT-as-a-judge as the most reliable metric to interpret.
@@ Appendix B Model Performance on Aurora-Bench with 5 Metrics
Copy PixInstruct GoT SE CM C-FT +Best-of-3 C-FDM +Best-of-3
@@
Table 6: Model performance at MagicBrush, Action-Genome, Something, WhatsUp and Kubric on Aurora-Bench. For C-FT and C-FDM We report both the model performance and their performance in the best-of-N distribution. We report the average GPT4o scores for each model at the bottom. We highlight the better GPT-4o scores for C-FT and C-FDM. We bold the best performance among all models, except Copy and best-of-N performances. SE: SmartEdit.
@@
Figure 6: Qualitative examples of the predicted next observation from the state-of-the-art specialised image editing models, and our models including C-FT and C-FDM, on Aurora-Bench.
Appendix C Qualitative Cases
In this section, we present additional qualitative examples from Aurora-Bench in Figure 6. We observe several common failure modes in image editing models. First, they sometimes fail to preserve the scene from the source observation (e.g., PixInstruct on Action-Genome and MagicBrush). Second, some models generate near-identical copies of the source as the target (e.g., GoT on Something-Something). Third, producing realistic outputs remains difficult, as seen in GoT’s result on Kubric. Finally, maintaining object consistency is also a challenge—SmartEdit alters the object in WhatsUp, and C-FDM does so in Something-Something.
Despite the challenges, we also observe several positive editing behaviours from C-FDM. On Action-Genome, C-FDM correctly predicts spatial changes, such as opening and closing a drawer, which requires a strong understanding of the spatial concepts. In Something-Something, it is the only model to accurately capture the spatial concept of “falling down.” On Kubric, it demonstrates basic counting ability by correctly adding one keyboard. In WhatsUp, C-FDM correctly grounds the action to the laptop, while other models mistakenly edit the monitor.
Appendix D Detailed Discussion for Chameleon’s Predicted Likelihoods
Figure 7: Comparison of predicted negative log-likelihoods (lower values indicate stronger model preference) for ground-truth real-world trajectories versus four types of negative trajectories. Top: Action prediction task for the IDP (observation $\times$ observation $\rightarrow$ action). Bottom: Next observation prediction task for the FDP (observation $\times$ action $\rightarrow$ observation). The legend shows the percentage of times the model prefers the ground-truth trajectory ($\uparrow$) over the negatives ($\downarrow$).
Figure 8: Comparison of negative log-likelihoods (lower values indicate stronger model preference) of the action predicted by IDM for ground-truth trajectories versus four types of negative trajectories.
From Figure 7, it emerges that Chameleon-7B displays a very limited preference for the ground-truth trajectories in a zero-shot setting. In the action prediction task (top panel), there is a slightly higher tendency to favour the ground-truth; however, even in the best case (counterfactual action), the model prefers the reference in only 58.1% of the samples. The high correlation in likelihoods indicates that the VLM struggles also on visual manipulations. In the next-observation prediction task (bottom panel), the VLM mostly fails in effectively differentiating the ground truth from the negatives. An exception to this is the copy manipulation, where the model can always tell them apart. Although the underlying reason remains uncertain, one plausible explanation for this behaviour is that the model’s ability to solve next-observation prediction tasks depends on their alignment with training sequences: for instance, it is plausible that Chameleon’s data rarely features two identical adjacent images. In Figure 8, we visualize the predicted likelihoods produced by Chameleon’s fine-tuned inverse dynamics model (IDM). We observe that fine-tuning on ground-truth trajectories substantially increases the model’s ability to distinguish ground-truth actions from negative alternatives. Specifically, the probability that the ground-truth action is assigned a higher likelihood increases from 55.6% to 73.2% under random-action negatives, and from 58.1% to 72.2% under counterfactual-action negatives. These results demonstrate the strong potential of learning effective inverse dynamics models directly from real-world trajectories. In summary, the zero-shot Chameleon-7B does not exhibit a preference for ground-truth trajectories over negative ones, constructed through action- or observation-based manipulations. However, it is possible to learn an effective IDM from the real-world trajectories.
Appendix E Details of Processing IDM Annotations for Unlabelled Videos
@@
Figure 9: Distributions of triplet log-likelihoods predicted by IDM on Movements-in-Time, UCF-101, and Kinetics-700, based on 7K synthetic triplets per dataset. Triplets are uniformly sampled from each action class while maximising overall predicted likelihoods.
@@
3:Initialize $S\leftarrow\emptyset$, and $\text{class\_counts}[c]\leftarrow 0$ for all $c\in\mathcal{C}$
@@
10:     $\text{class\_counts}[c]\leftarrow\text{class\_counts}[c]+1$
@@
We present the raw dataset statistics before sampling for Movements-in-Time, UCF-101 and Kinetics700 in Table 7. Figure 9 shows the distribution of IDM’s predicted scores across action classes in Movements-in-Time, Kinetics700, and UCF-101. The predicted likelihoods are nearly uniform within each class, indicating that our sampling method maintains both class diversity and high overall likelihoods. The sampling procedure for IDM-annotated trajectories is detailed in Algorithm 1.
@@
Table 7: Dataset statistics for the video and triplets from the trajectories annotated by IDM. OPV: observations (i.e., extracted key-frames) per video, APV: actions per video, WPA: words per action.
Appendix F Prompt Template for Using GPT4o-as-a-Judge
Prompt Template for GPT4o-as-a-judge Evaluation You are a professional digital artist. You will have to evaluate the effectiveness of the AI-generated image(s) based on the given rules. You will have to give your output in a valid way of a Python dictionary format (Keep your reasoning concise and short.): {{"score": [...], "reasoning": "..." }} and don’t output anything else. Two images will be provided: The first being the original AI-generated image The second being an edited version of the first. The objective is to evaluate how successfully the editing instruction has been executed in the second image. Note that sometimes the two images might look identical due to a failure in image editing. From a scale of 0 to 10: A score from 0 to 10 will be given based on the success of the editing. A second score from 0 to 10 will rate the degree of minimal editing. Editing instruction: {instruction}
Figure 10: Prompt template used for GPT-4o-as-a-judge evaluation.
We provide the prompts used for evaluating image editing performance with GPT-4o in Figure 10. We use GPT-4o-2024-11-20. The final score is the average of the minimum value of the two scores for each sample, as in (Fang et al., 2025).
Appendix G Detailed GPT4o Scores for Editing Success and Minimal Editing
(a) Detailed GPT4o scores for C-FDM trained with the standard loss.
(b) Detailed GPT4o scores for C-FDM trained with the L2-weighted loss.
Figure 11: GPT4o scores’ distributions of editing success (ES) and minimal editing (OE) for C-FDM trained with standard loss or our loss-weighting method.
@@ Appendix G Detailed GPT4o Scores for Editing Success and Minimal Editing
Appendix H Implementation Details
H.1 Chameleon Dynamics Model
@@
We fine-tune the Chameleon-7B checkpoint from the Anole-7B version (Chern et al., 2024) to predict the action given a pair of observations, framed as an action-prediction task. The model is trained on a merged dataset from Action-Genome, Kubric, MagicBrush, Something-Something from AURORA’s annotated trajectories, and 15K EPIC-Kitchens processed by us. We downsample Kubric’s trajectories to 10K. Training is performed for 10 epochs with a batch size of 64, using a learning rate of 2e-4 and cosine scheduling (500 warm-up steps). We use bfloat16 mixed-precision training and apply LoRA (Hu et al., 2022) for parameter-efficient fine-tuning (rank 16, $\alpha=32$, dropout 0.05). Only the completion loss is used to optimise the generation of action. Training is conducted on 4 NVIDIA-H100-80GB-HBM3 GPUs using DeepSpeed for distributed optimisation.
H.2 C-FT Baseline
@@
We fine-tune the Chameleon-7B checkpoint from the Anole-7B version (Chern et al., 2024). The model is trained on a combined dataset from Action-Genome, Kubric, MagicBrush, and Something-Something, formatted as the image editing task. We downsample Kubric’s trajectories to 10K. Training is conducted for 40 epochs with a batch size of 96 using the AdamW optimiser and a cosine learning rate scheduler (learning rate of 5e-4, 400 warm-up steps). We use mixed-precision training with bfloat16 and apply LoRA (Hu et al., 2022) for efficient fine-tuning (rank 16, $\alpha=32$, dropout 0.05). We only train the model with the truncated loss from the completion part. We use 4 NVIDIA-H100-80GB-HBM3 GPUs with DeepSpeed for distributed training. During inference, we apply a logits processor to mask out non-image tokens, set the temperature to 1, and use top-1 sampling. We observe that temperature is critical in controlling model behaviour: lower values often cause the model to copy the source observation instead of generating meaningful edits.
H.3 Chameleon FDM
@@
We fine-tune the Chameleon-7B checkpoint from the Anole-7B version (Chern et al., 2024). The model is trained on a combined dataset from Action-Genome, Kubric, MagicBrush, Something-Something from AURORA’s annotated trajectories, together with 7K trajectories from Movements-in-Time, 7K trajectories from UCF-101 and 7K trajectories from Kinetics700, formatted as the image editing task. Again, we downsample Kubric’s trajectories to 10K. Training is conducted for 40 epochs with a batch size of 96 using the AdamW optimiser and a cosine learning rate scheduler (learning rate of 5e-4, 400 warm-up steps). We use mixed-precision training with bfloat16 and apply LoRA (Hu et al., 2022) for efficient fine-tuning (rank 16, $\alpha=32$, dropout 0.05). We only train the model with the truncated loss from the completion part, but we weight the image tokens using L2 strategy as introduced in Section 3. We use 4 NVIDIA-H100-80GB-HBM3 GPUs with DeepSpeed for distributed training. We use the same hyperparameters as C-FT during the inference time.
H.4 Computing Resources
@@
All training experiments were conducted on a compute node equipped with 4× NVIDIA H100 80GB GPUs, 256 CPU cores, and 256GB of memory. The total GPU hours required for training C-FT, C-FDM, and IDM were approximately 200, 400, and 100 hours, respectively.
@@
For inference, we used a single NVIDIA A100 80GB GPU with 8 CPU cores and 128GB memory. Inference for C-FT and C-FDM takes approximately 1 GPU hour per model. When applying verification with $K=8$, inference time increases to around 8 GPU hours. IDM only takes around 0.3 GPU hours for inference.
H.5 Assets and Licenses
@@
Appendix I Details of Human Evaluation
@@
Appendix J Safeguards
C-FDM performs observation prediction through image generation and, while its outputs are task-specific, we acknowledge that any generative model may carry potential for misuse. To mitigate these risks, we commit to the following safeguards upon release:
@@
Appendix K LLMs Usage Declaration
We declare that the large language model (LLM) was only used to assist in minor tasks, including revising the manuscript for grammatical correctness, improving phrasing, and performing small technical implementations such as debugging code snippets. All scientific ideas, results, analyses, and conclusions presented in this paper are entirely the work of the authors.